some students simplify 2x^2-32/x-4 as 2x-8/x-4 =2x-8. choose values of x to mshow this is not true. Then simplify it correctly.
Answer by stanbon(75887)

The wrong way:
2x^2-32/x-4 as 2x-8/x-4 =2x-8
Example:
Let x = 2
You would get 2*4-32 /2-4 = 2*2-8
8-32/-2 = 4-8
-24/-2 = -4
12 = -4
Which is wrong.
==============================
The Right way:
(2x^2-32)/(x-4)
------
= [2(x^2-16)]/(x-4)
= [2(x-4)(x+4)]/(x-4)
= 2(x+4)
==============
Example: Let x = 2
You get
(2*4-32)/(2-4) = 2(2+4)
(8-32)/-2 = 2(6)
-24/-2 = 12
12 = 12
Which is correct.
=====================
Cheers,
Stan H.